In order to inc the resistance of  a  given wire of uniform cross section to four times it's value, a fraction of length of the wire becomes 3/2 times the original length. Show that the value of this fraction is 1/8.


Grade:12

1 Answers

Pratham Ashish
14 years ago
Hi, tushar der is smtng wrong with ur question. Let us consider that v increase the length of the whole wire to 3/2 times of it original length , still the R wud become 9/4 R.IT can't be 4R in any case. Consider length is increase to 3/2 times,then its cross section area will become 2/3 times of its original, the resistance is proportional to length and inversly proportional to area , hence the new resistance wud b {(3/2L)/(2/3A)}=9/4*L/A = 2.25R. So i request u to plz confirm ur qn & repost ur doubt.